OpenStudy (anonymous):

Using the domain values {-1, 0, 4}, find the range values for relation y=2x-10. Range={-11, -10, -6} Range={12, -10, 2} Range={-12, -10, -2} None of the above

OpenStudy (anonymous):

substitute one of the domain value in place of x in the equation.... find y example for domain value -1 sub '-1' in place of x y=2(x)-10 y=2(-1)-10 y=-2-10 y=-12 do this for all the others to find different values of y...... the set containing the values is the range...

OpenStudy (anonymous):

oh alright but its the last one ?

OpenStudy (anonymous):

yep!!

OpenStudy (anonymous):

if i i work one out can you tell me if im right

OpenStudy (anonymous):

sure!